### v1

# Multi-Layer Cycle Benchmarking for high-accuracy error characterization

Abstract
Accurate noise characterization is essential for reliable quantum computation. Effective Pauli noise models have emerged as powerful tools, offering detailed description of the error processes with a manageable number of parameters, which guarantees the scalability of the characterization procedure. However, a fundamental limitation in the learnability of Pauli fidelities impedes full high-accuracy characterization of both general and effective Pauli noise, thereby restricting e.g., the performance of noise-aware error mitigation techniques. We introduce Multi-Layer Cycle Benchmarking (MLCB), an enhanced characterization protocol that improves the learnability associated with effective Pauli noise models by jointly analyzing multiple layers of Clifford gates. We show a simple experimental implementation and demonstrate that, in realistic scenarios, MLCB can reduce unlearnable noise degrees of freedom by up to $75\%$, improving the accuracy of sparse Pauli-Lindblad noise models and boosting the performance of error mitigation techniques like probabilistic error cancellation. Our results highlight MLCB as a scalable, practical tool for precise noise characterization and improved quantum computation.
I Introduction
A wide variety of characterization techniques, serving diverse purposes and spanning different ranges of applicability, have been developed to date. For instance, Gate Set Tomography (GST) provides a comprehensive method to determine individual errors associated with executing sets of quantum operations [1], but its exponential scaling in sampling complexity limits its application to large systems [2]. On the other hand, holistic techniques like Cross-Entropy Benchmarking (XEB) [3, 4] or scalable variants of Randomized Benchmarking (RB) [5, 6, 7] aim to benchmark large Quantum Processing Units (QPUs) by assessing their average quality.
Recently, considerable attention has been devoted to characterization schemes that lie between these extremes, offering greater detail than holistic techniques while still remaining scalable [8, 9, 10, 11, 12]. This trend is also driven by the need to implement noise-aware error mitigation techniques [13, 8, 14, 15], which require detailed knowledge of the noise associated with the parallel execution of instructions on large QPUs. Achieving this without incurring the prohibitively high sampling complexity of GST typically requires implementing averaging techniques to recast the original noise into more manageable forms. Equally important is the development of simpler, effective noise models that capture the most relevant features of the noise on the whole QPU while featuring a manageable number of parameters [8, 16, 11].
A prominent example of noise recasting techniques is Randomized Compiling (RC) applied to Clifford gates, which can approximately average, or twirl, noise into so-called Pauli noise [17, 18]. The latter is a type of stochastic noise modeled by Pauli channels, where only errors associated to any Pauli operator occur with a given probability. Incidentally, this makes RC also a valuable error suppression tool, as it gets rid of a large portion of detrimental, so-called coherent errors that get twirled into less detrimental, incoherent ones. This highlights the appeal and practical significance of characterizing Pauli noise [8, 16]. Its relevance is further amplified by the possibility to develop and use effective and scalable versions of Pauli noise, where reasonable assumptions about noise locality can significantly reduce the number of parameters to be determined [8, 9, 16].
Despite these nice features, the characterization of gate-dependent Pauli noise remains fundamentally limited by Pauli learnability constraints, as demonstrated in Ref [19] and confirmed experimentally [8, 20, 17]. As in GST [2, 21], gauge freedom introduces ambiguities, making it impossible to determine certain parameters of Pauli channels independently from State Preparation and Measurement (SPAM) errors. Specifically, Ref. [19] shows that for a set of Clifford gates, the number of these unlearnable noise Degrees of Freedom (DOF) in Pauli channels scales exponentially with the number of qubits $n$. In this context, Cycle Benchmarking (CB) has emerged as an convenient method for accurately determining all learnable DOF of Pauli channels [19].
In this work, we specifically consider the characterization of scalable and effective Pauli noise models and present a enhanced version of CB called Multi-Layer Cycle Benchmarking (MLCB), which significantly increases the number of learnable DOF by leveraging the constraints inherent in the effective noise model. Unlike standard CB implementations [8, 16], MLCB simultaneously analyzes multiple sets of parallel Clifford gates, rather than characterizing each set in isolation. After providing a general framework for implementing MLCB, we focus on the characterization of the IQM GarnetTM [22] quantum processor, which features $20$ superconducting qubits in a square topology. Our results show that MLCB can achieve a remarkable $75\%$ reduction in the number of unlearnable DOF in practically relevant scenarios. Moreover, we experimentally highlight how MLCB-enabled learnable DOF differ significantly from estimates derived through conventional low-accuracy approaches. Additionally, we numerically show how this enhanced learnability offered by MLCB translates into more accurate characterization of effective noise models, focusing on the Sparse Pauli-Lindblad (SPL) noise model introduced in Ref. [8] and widely used to inform noise-aware error mitigation strategies [8, 14, 15]. In this respect, we also investigate numerically how MLCB reduces residual biases in Probabilistic Error Cancellation (PEC), a prominent error mitigation technique that requires precise noise characterization to achieve optimal results [13, 8, 23].
The rest of the paper is organized as follows: Section II provides an overview of the CB technique for Clifford gate layer characterization and discusses Pauli noise learnability constraints. Section III introduces scalable Pauli noise models, with a focus on the SPL model. The MLCB framework is presented in Section IV, followed by its experimental testing on the IQM GarnetTM platform in Section V. Finally, Sections VI and VII illustrate how MLCB enhances characterization accuracy and improves the performance of noise-aware error mitigation techniques.
II Characterization of a Clifford layer
In general, a quantum circuit consists of a list of instructions that can be executed by the quantum hardware, such as quantum gates, reset or state preparation and measurements. The set of qubits targeted by a specific instruction is known as the instruction’s support. We define a layer as a set of instructions featuring non-overlapping supports, which is executed in parallel (although not necessarily simultaneously)  [^1] In the literature, a layer is sometimes also referred to as a cycle [16]. This allows us to conveniently express a generic circuit in terms of a sequence of layers, executed at given time steps. In practically relevant scenarios, the support of native instructions often consists of either one or two qubits. Entangling two-qubit gates are typically Clifford operations, such as iSWAP, CNOT and CZ (the latter being native on IQM’s hardware). When executed on the QPU, two-qubit (2Q) gates are generally noisier than general single-qubit (1Q) rotations. Moreover, because of spatial crosstalk, the execution of a given gate can affect the behavior of neighbouring qubits. Characterizing the noise associated with the execution of a whole Clifford layer $C$, consisting of several 2Q gates, is thus crucial for assessing the errors impacting practical applications. In the following, we introduce the necessary notation by defining the concepts of Pauli twirling, Pauli maps, and the associated Pauli fidelities. We then provide a concise overview of the standard CB protocol [25], which enables the high-accuracy determination of all learnable Pauli fidelities associated with general Pauli noise [19]. This protocol serves as the foundation for the development of our proposed MLCB technique.
II.1 Determining the Pauli fidelities
The ideal implementation of $C$ on a $n$-qubit QPU is represented by a unitary operator $U_{C}\in\mathbb{U}(2^{n})$, corresponding to the ideal quantum channel $\phi(U_{C})$, defined as $\phi(U_{C})[\rho]:=U_{C}\rho U_{C}^{\dagger}$. However, in practice, assuming noise being context-independent and probability-preserving, the actual execution of the layer can be modelled as $\mu(U_{C}):=\phi(U_{C})\circ\tilde{\Lambda}_{C}$, where the completely positive and trace-preserving (CPTP) map $\tilde{\Lambda}_{C}$ captures the effects of the noise. To simplify the form of the noise channel $\tilde{\Lambda}_{C}$ and enable efficient noise characterization, a key technique is twirling. Formally, the twirling of the map is performed by introducing additional layers, before and after the noisy Clifford layer $\phi(U_{C})$, consisting of 1Q operators randomly sampled from the twirling set $\mathbb{T}$. Unless explicitly stated otherwise, in the following we will focus on the standard Pauli twirling, with $\mathbb{T}$ consisting of all possible product combination of 1Q Pauli operators, i.e. $\mathbb{T}=\mathbb{P}_{n}:=\{I,X,Y,Z\}^{\otimes n}$. Instead of executing $U_{C}$, we first implement a twirling layer $T\in\mathbb{T}$, followed by $U_{C}$ and then by the conjugated layer $T_{C}:=U_{C}TU_{C}^{\dagger}\in\mathbb{T}$, which also belongs to $\mathbb{T}$ since the Pauli group is preserved under Clifford conjugation. If we assume that the noise map $\Omega$ associated with the execution of a layer $T$ of 1Q gates is gate-independent, i.e. $\nu(T):=\Omega\circ\phi(T)$ $\forall T$, one can show that uniformly sampling $T$ from the twirling set $\mathbb{T}$ and subsequently averaging the results, denoted as $\langle\cdot\rangle_{T\in\mathbb{T}}$, leads to the following effective evolution
$\begin{split}&\langle\nu(T_{C})\circ\mu(U_{C})\circ\nu(T)\rangle_{T\in\mathbb{T}}\\ &\qquad=\Omega\circ\phi(U_{C})\circ\langle\phi(T)\circ\tilde{\Lambda}_{C}\circ\Omega\circ\phi(T)\rangle_{T\in\mathbb{T}}\\ &\qquad=\Omega\circ\phi(U_{C})\circ\Lambda_{C},\end{split}$ (1)
where
$\Lambda_{C}[\rho]=\sum_{\alpha}p_{\alpha}P_{\alpha}\rho P_{\alpha}^{\dagger}\quad(P_{\alpha}\in\mathbb{P}_{n}),$ (2)
is an effective Pauli channel associated with the execution of the Pauli-dressed Clifford layer $C$. The channel $\Lambda_{C}$ corresponds to the noise of the original Clifford layer $C$, which contributes the majority of the noise, together with the noise of the twirling layer $T$. Note that, even if the noise associated with the latter has some gate-dependent components, the whole construction still holds up to quadratic corrections in the deviations from the gate-independent assumption [16].
We can now focus on characterizing the Pauli map $\Lambda_{C}$: since it is a linear map, it is straightforward to see that every Pauli operator $P_{\beta}\in\mathbb{P}_{n}$ is an eigenvector of the map, i.e., $\Lambda_{C}[P_{\beta}]=f_{\beta}P_{\beta}$ $\forall P_{\beta}\in\mathbb{P}_{n}$, with the associated eigenvalue $f_{\beta}$. Accurately characterizing the eigenvalues $f_{\beta}$, also known as Pauli fidelities, is one of the main objectives of this paper. Indeed, a perfect knowledge of all the Pauli fidelities $\{f_{i}\}$ allows to determine the elements $\{p_{i}\}$ of the probability distribution in Eq. (2) via the Walsh-Hadamard transform $p_{\alpha}=2^{-n}\sum_{\beta}f_{\beta}(-1)^{\langle\alpha,\beta\rangle_{sp}}$, with $\langle\alpha,\beta\rangle_{sp}$ being the symplectic inner product of Pauli operators $P_{\alpha}$ and $P_{\beta}$, which is zero if they commute and one otherwise.
If it was possible to implement the noise map $\Lambda_{C}$ without the associated unitary $\phi(U_{C})$, the generic Pauli fidelity $f_{\alpha}$ could be measured by implementing a simplified version of the CB protocol, consisting of the following steps: (i) prepare a $+1$ eigenstate of the operator $P_{\alpha}$, (ii) apply $\Lambda_{C}$ a number of times $d$ and, (iii) measure the expectation value of $P_{\alpha}$. Indeed, for $d=1$, one obtains
$\begin{split}&\text{Tr}\left\{P_{\alpha}\Lambda_{C}[\rho_{\alpha}]\right\}=\frac{\text{Tr}\left\{P_{\alpha}\Lambda_{C}[P_{\alpha}]\right\}}{2^{n}}=f_{\alpha}.\end{split}$ (3)
For generic $d$, in the presence of SPAM errors and by using readout twirling [26, 27], one can show that the expectation value is given by $A_{\alpha}(f_{\alpha})^{d}$  [^2] Without readout twirling one would generally get a constant shift $A_{\alpha}(f_{\alpha})^{d}+B_{\alpha}$ By performing this procedure for different values of $d$ and fitting the results to an exponential curve, one can determine $f_{\alpha}$ with high accuracy. Here, high accuracy signifies that the quantity can be measured with multiplicative precision in a SPAM-robust manner, as the SPAM coefficients $A_{\alpha}$ are independent of $d$ [25].
However, in practice, it is not feasible to apply the map $\Lambda_{C}$ in isolation, as the entire entire dressed Clifford layer $C$ needs to be implemented. The resulting characterization protocol becomes what is known as standard CB and the learnability of the Pauli map $\Lambda_{C}$ becomes limited, as we will discuss in the following.
Figure 1: Illustration of the standard CB protocol. In the state preparation stage (left brown area) an eigenstate of a Pauli operator is prepared. The Clifford layer $C$ subject to characterization (in this case, consisting of several parallel CZ gates shown in blue) is then implemented $d$ times, alternating it with layers of 1Q gates (white boxes) which are needed to implement RC. This can be interpreted as the implementation of $d$ dressed Clifford layers. After a final layer of 1Q gates which ensures that RC does not modify the unitary associated with the circuit, the measurement stage (right brown box), which also incorporate readout-twirling, provide us with the expectation value of a Pauli operator (typically the same operator as considered in the state preparation stage).
II.2 Limited learnability of Pauli noise
The standard CB protocol is illustrated in Fig. 1 where, after a state preparation stage (left brown region), the twirled layer $C$ is implemented $d$ times and followed by a final measurement stage (right brown region). Since adjacent 1Q layers can be compiled together [18], the central region of the circuit consists of $d$ instances of $C$ (consisting of parallel 2Q gates shown in blue) which alternate with $d+1$ 1Q gate layers (white boxes). By preparing a $+1$ eigenstate of a Pauli operator $P_{\alpha}$ and measuring its expectation value (using readout twirling), the measurement outcome $\mathsf{O}_{d}^{\alpha}$ reads
$\mathsf{O}_{d}^{\alpha}\propto 2^{-n}\;\text{Tr}\left\{P_{\alpha}\,\Omega\circ\left(\phi(U_{c})\circ\Lambda_{C}\right)^{\circ d}[P_{\alpha}]\right\}.$ (4)
The extra noise map $\Omega$, originating from the last “unpaired” layer of 1Q gates, is not problematic since it can be effectively considered as part of the SPAM error. By contrast, the presence of $\phi(U_{c})$ adds one important complication with respect to the simple scenario discussed before. Indeed, while the Clifford nature of $C$ assures that a Pauli operator is mapped to another Pauli $\phi(U_{C})[P_{\alpha}]=U_{C}P_{\alpha}U_{C}^{\dagger}=P_{\beta}$, there is no guarantee that $P_{\alpha}=P_{\beta}$. This brings us to the definition of the orbit $\mathcal{R}^{C}_{\alpha}$ of a Pauli operator $P_{\alpha}$ as the set of unique Pauli operators that can be reached via the repeated application of the Clifford layer $C$, i.e. $\mathcal{R}^{C}_{\alpha}=\{U_{C}^{j}P_{\alpha}(U_{C}^{\dagger})^{j}\}|j\in\mathbb{N}\}$ [16]. Note that the cardinality of each orbit is upper bounded by $|\mathcal{R}^{C}_{\alpha}|\leq m$, where $m$ is the smallest positive integer such that $(U_{C})^{m}=I$, i.e., the order of Clifford group element $U_{C}$. If $|\mathcal{R}^{C}_{\alpha}|=1$, then the CB experiment described in Eq. (4) results in $\mathsf{O}_{d}^{\alpha}=A_{\alpha}(f_{\alpha})^{d}$. It is thus possible to determine the individual fidelity $f_{\alpha}$ with high accuracy. However, if $|\mathcal{R}^{C}_{\alpha}|>1$, only the product of the corresponding fidelities can be accessed with high accuracy. Indeed, considering $d=k|\mathcal{R}^{C}_{\alpha}|$ (with $k\in\mathbb{N}$), the resulting measurement outcome of the CB experiment is given by the product of the Pauli fidelities, or more explicitly
$\mathsf{O}_{d}^{\alpha}=A_{\alpha}\left(\prod_{\beta\in\{\gamma|P_{\gamma}\in\mathcal{R}^{C}_{\alpha}\}}f_{\beta}\right)^{k}.$ (5)
Importantly, 1Q Clifford gates, such as $S=\sqrt{Z}$ and $\sqrt{X}$, can map each of the three 1Q Pauli operators ($X$, $Y$, and $Z$) into either of the other two. As a result, if a given orbit contains different Pauli operators that share same support, its cardinality can be reduced, thus enabling the determination of more quantities with high accuracy. The combination of this approach, known as interleaved CB, with the standard CB protocol allows to access all the learnable DOF of general Pauli noise [19]. By contrast, the cardinality of orbits consisting of operators with different support cannot be reduced further. Their existence is thus associated with the presence of unlearnable DOF, which can be also understood as a consequence of a fundamental gauge freedom, similar to the one encountered in the context of GST [19, 2]. Specifically, out of the $14^{n}-$ 1 DOF of general Pauli noise, the fundamentally unlearnable ones amount to $2^{n-c}$, where where $c$ is the number of connected components of the pattern transfer graph obtained from the analysis of the orbits (see Ref. [19] for more details).
Table 1: Conjugation of non-trivial Pauli operators in $\mathbb{P}_{2}$ under a single CZ (described by unitary $U_{C}$) and for a single CZ with extra S gates (unitary $\bar{U}_{C}$). Conjugated Pauli strings in bold are different from the original $P_{\alpha}$.
$P_{\alpha}$ IX IY IZ XI XX XY XZ YI YX YY YZ ZI ZX ZY ZZ
$U_{C}P_{\alpha}U_{C}^{\dagger}$ ZX ZY IZ XZ YY YX XI YZ XY XX YI ZI IX IY ZZ
$\bar{U}_{C}P_{\alpha}\bar{U}_{C}^{\dagger}$ ZY ZX IZ YZ XX XY YI XZ YX YY XI ZI IY IX ZZ
For the sake of concreteness, it is convenient to discuss a simple example: the case of a single CZ gate performed on a $2$-qubit device. In this case, $U_{C}^{2}=I$ limits the cardinality of the orbits to $2$, as it emerges also from Tab. 1, which shows the conjugation of all non-trivial Pauli operators in $\mathbb{P}_{2}$ under a single CZ . Allowing for the usage of additional interleaved $S$ gates, which map $X$ into $Y$ and vice versa, there are $7$ Pauli fidelities which are directly learnable (as they are associated with orbits with cardinality $1$). The remaining $8$ fidelities are not directly measurable but the CB protocols allow to determine, with high accuracy, $8$ products of fidelity pairs. The latter are not all independent from each other since, for example, $f_{IY}f_{ZX}=(f_{IY}f_{ZY})(f_{IX}f_{ZX})/(f_{IX}f_{ZY})$. Still, they allow to impose a total of $6$ high-accuracy constraints on the parameters of the noise model, leaving us with only $2$ unlearnable DOF, which can be identified, for example, with fidelities $f_{IX}$ and $f_{XI}$ (this identification is clearly not unique). More details concerning this simple example are provided in Appendix A.
We conclude this section by emphasizing that the limited learnability of Pauli noise—stemming from the inherent challenge of independently characterizing state preparation and measurement errors [19]—also impacts alternative protocols developed to measure Pauli fidelities. Examples include tensor-network-based noise learning [11] and Averaged Circuit Eigenvalue Sampling (ACES) [9, 12, 10]. For instance, the ACES protocol can reconstruct all relevant fidelities associated with a local Pauli noise model only under the assumption of perfect state preparation [9, 12].
III Scalable Pauli characterization
The Sparse Pauli-Lindblad (SPL) model, introduced in Ref. [8], provides a scalable framework for characterizing Pauli noise. This model has already been implemented on a 127-qubit device, yielding valuable characterization data that allowed the implementation of noise-aware error mitigation protocols across the entire QPU [15]. Another scalable characterization scheme, known as Cycle Error Reconstruction (CER) [16], has similarly been leveraged to inform error mitigation strategies [20]. Both methods achieve scalability by relying on the reasonable assumption that the most relevant errors have a small, local support, reflecting the QPU’s topology. While our findings are beneficial for both approaches, this work primarly focuses on the SPL model, whose key features are outlined in the remainder of this section.
III.1 The SPL model
The SPL model assumes that $\Lambda_{C}$, the effective Pauli noise map in Eq. (2) associated with the dressed Clifford layer $C$, is generated by a Lindbladian whose jump operators are local and low-weight Pauli operators [8]. Specifically, it can be expressed as
$\Lambda_{C}(\rho)=e^{\mathcal{L}}(\rho)=\exp\sum_{\alpha\in\mathcal{K}}\lambda_{\alpha}^{C}(P_{\alpha}\rho P_{\alpha}-\rho),$ (6)
where the set $\mathcal{K}$ consists of Pauli strings with a support involving up to $w$ neighbouring qubits (with respect to the QPU’s topology). To achieve a full characterization of the error model, all the model coefficients $\lambda^{C}_{\alpha}$ need to be determined. Considering a maximum weight $w=2$ — a common assumption in the literature [8, 15, 29] and which we adopt throughout this paper — the number of model parameters associated with $\Lambda_{C}$ is given by $|\mathcal{K}|=3n+9p$. Here, $n$ is the number of qubits and $p$ is the number of nearest-neighbour pairs. For a square topology, at scale, we have $p\sim 2n$ which leads to the linear scaling $|\mathcal{K}|\sim 21n$.
The model parameters $\lambda_{\alpha}^{C}$ cannot be accessed directly but they are related to the Pauli fidelities, associated with the dressed Clifford layer $C$, via
$\log(f_{\alpha})=-2\sum_{\beta\in\mathcal{K}}M_{\alpha\beta}\lambda_{\beta},$ (7)
where $M_{\alpha\beta}=\langle\alpha,\beta\rangle_{sp}$. For brevity, we omitted the superscript $C$ in Eq. (7) as well as in the remainder of the manuscript. It is convenient to generalize Eq. (7) to take into account that, in some cases, only products of Pauli fidelities are known with high accuracy. To this end, we define the full $|\mathcal{K}|$-dimensional vector of model parameters $\vec{\lambda}$  [^3] Implicitly this means that we define an ordering of the set $\mathcal{K}$, which maps a Pauli string $\alpha$ to an integer $i\in\mathbb{N}$ so that the $i$-th element of $\vec{\lambda}$ (denoted by $\lambda_{i}$) is given by $\lambda^{C}_{\alpha}$.. Similarly, we consider an $F$-dimensional vector $\vec{\xi}$ whose $j$-th element consists of a given product of fidelities, associated with the Pauli strings in a given set $\mathcal{X}_{j}$ (which typically corresponds to some orbit), i.e. $\xi_{j}=\prod_{\alpha\in\mathcal{X}_{j}}f_{\alpha}$. Then we can generalize Eq. (7) to
$\log(\vec{\xi})=-2\bar{M}\vec{\lambda}$ (8)
where the logarithm is applied element-wise and $\bar{M}$ is a rectangular $F\times|\mathcal{K}|$ matrix whose $j$-th row reads $\bar{M}_{jk}=\sum_{\alpha\in\mathcal{X}_{j}}M_{\alpha k}$. If enough Pauli fidelities (or products thereof) are known such that the rank of $M$ is equal to $|\mathcal{K}|$, then Eq. (8) can be inverted and the model parameters determined, typically via a non-negative least square fit
$\vec{\lambda}(\vec{\xi})={\rm argmin}_{\lambda_{k}\geq 0}\;||\bar{M}\vec{\lambda}+\log(\vec{\xi})/2||_{2}^{2}.$ (9)
Once all the model parameters $\vec{\lambda}$ are determined, the SPL model proves to be very useful. For example, it can be readily used to effectively amplify or reduce the circuit noise, enabling the execution of error mitigation techniques such as PEC [13, 8], tensor-network error mitigation (TEM) [14] and zero noise extrapolation [15, 31] with probabilistic amplification. Indeed, one can show that the (quasi)probabilistic implementation of the map  [^4] This can be done by randomly sampling circuits containing Pauli errors $P_{k}$ according to the (quasi)probability distribution defined by coefficients $w_{k}^{(\beta)}$ [8]
$\Lambda_{C}^{(\beta)}(\rho)=\prod_{k\in\mathcal{K}}\left(w_{k}^{(\beta)}\rho+(1-w_{k}^{(\beta)})P_{k}\rho P_{k}\right)$ (10)
before the layer $C$ effectively scales the overall noise as $\Lambda_{C}\circ\Lambda_{C}^{(\beta)}=e^{\beta\mathcal{L}}$, provided that $2w_{k}^{(\beta)}=1-\exp[-2(\beta-1)\lambda_{k}]$ [33].
III.2 Effects of limited learnability
Let us consider the characterization of a Clifford layer $C$ composed of parallel CZ gates - a scenario of significant practical relevance [22, 34, 35]. In order to fit the corresponding SPL model, it is sufficient to know [8] (i) all the $3n$ fidelities associated with unit-weight Pauli operators and (ii) all products of fidelity pairs $f_{\alpha}f_{\beta}$ where $\alpha\in\mathcal{K}$ is a weight two Pauli string and $\beta$ is the Pauli string associated to the conjugation of $P_{\alpha}$ under the layer $C$, possibly conjugated with a second Clifford layer, consisting only of 1Q gates  [^5] Indeed, it is required that each one of the two elements of $\beta$ in the support of $\alpha$ is either the identity or equal to the corresponding element of $\alpha$. This can be always ensured by composing the layer $C$ with a second Clifford layer, consisting only of 1Q gates, in the same spirit as the interleaved CB [8]. While the combination of standard and interleaved CB provides us with high-accuracy values for all the fidelity products required in (ii), condition (i) is highly impacted by the limited learnability of Pauli noise and some fidelities will need to be estimated using different, less accurate approaches.
In the previous section, we demonstrated that a single, isolated CZ gate has two unlearnable DOF, which can be conveniently associated with the unit-weight fidelities $f_{IX}$ and $f_{XI}$. Leveraging this fact, along with condition (i), we can immediately compute the number of unlearnable DOF for an SPL characterization of a layer containing $n_{g}$ parallel CZ gates, which totals $2n_{g}$. These can be also conveniently associated with $2n_{g}$ unlearnable unit-weight fidelities, each one of them featuring a single $X$ operator on one qubit belonging to the layer’s support.
Various approaches have been proposed to estimate these unlearnable fidelities. One option involves performing additional experiments using unit-depth ($d=1$) variants of the CB protocols, where the layer to be characterized is applied only once. However, in this case, it is not possible to distinguish the Pauli fidelities of interest from SPAM errors, which can be effectively removed only by assuming perfect state preparation [8, 19, 33]. As a result, the data obtained from this method lacks high accuracy. Another approach assumes symmetries among the Pauli fidelities. For example, it may be assumed that certain unlearnable fidelities are equal to their conjugates under the action of $C$. In the case of a single CZ gate, this corresponds to assuming, for instance, $f_{XI}=f_{XZ}$ and then estimate the individual unlearnable fidelity as $f_{XI}=\sqrt{f_{XI}f_{XZ}}$, where the product on the right hand side is learnable. If this symmetry assumption holds, the SPL model can be fit using only high-accuracy data. However, if the assumption is invalid, this method introduces errors.
Regardless of the chosen strategy for addressing limited learnability, once conditions (i) and (ii) are satisfied, all information about the (products of) fidelities can be encapsulated in the vector $\vec{\xi}$. It is useful to represent $\vec{\xi}$ as a concatenation $\vec{\xi}=(\vec{\Xi},\vec{\zeta})$, where $\vec{\Xi}$ contains only learnable (products of) fidelities, and $\vec{\zeta}$ comprises unlearnable ones. At this stage, the full-rank matrix $\bar{M}$ can be constructed, enabling the SPL model parameters $\vec{\lambda}$ to be determined by inverting Eq. (8). As shown in Refs. [19, 8] and further discussed in Sec.VI, the limited accuracy associated with $\vec{\zeta}$ impacts the overall precision of the scalable error characterization.
Figure 2: A QPU with square topology and four layers of parallel CZ gates, identified with blue, green, red and orange colors, that covers all nearest-neighbour connections between the qubits.
So far, we have focused on the characterization of a single layer $C$. However, in realistic scenarios, multiple distinct layers must often be addressed. A particularly relevant example is depicted in Fig. 2, which illustrates a large $n$-qubit QPU with a square topology and four layers of parallel CZ gates, color-coded as blue (B), green (G), red (R), and orange (O). These four layers collectively cover all nearest-neighbor connections on the QPU. Combined with arbitrary 1Q gate layers, they enable the construction of non-trivial circuits that facilitate rapid entanglement distribution across the whole QPU [37]. Such dense configurations are useful for a variety of applications, including simulating the dynamics of lattice models [15] and implementing variational quantum algorithms [38].
In the limit of large $n$, the number of CZ gates per layer scales linearly as $n_{g}\sim 2n$, as does the number of SPL parameters per layer, $|\mathcal{K}|\sim 21n$. Characterizing all four layers requires determining a total of $\sim 84n$ parameters. If each layer is characterized independently, the total number of unlearnable quantities amounts to $2n_{g}\sim 4n$. Unlike the (non-scalable) full characterization of Pauli noise—where the unlearnable DOF constitute an exponentially small fraction of the learnable ones [19]—this example shows that for large $n$ and considering the SPL model, the ratio of unlearnable DOF to total parameters approaches a constant value of around $4/84\sim 4.8\%$. As we will discuss in Sections VI and VII, this small fraction is enough to significantly hinder the overall quality of error characterization and the effectiveness of noise-aware error mitigation techniques.
III.3 Number of required experiments
Compared to strategies based on depth-one circuits, approaches relying on symmetry assumptions have the significant advantage of not requiring additional experiments. In these cases, the entire experimental runtime is indeed allocated to executing standard and interleaved CB protocols. For the sake of concreteness, we still focus on QPU s with a square lattice topology, as illustrated in Fig. 2. In this scenario, it has been shown that the (products of) fidelities associated with all orbits $\mathcal{R}^{C}_{\alpha}$ for $\alpha\in|\mathcal{K}|$ can be measured using standard CB protocols with just nine distinct state preparation and measurement stages [8]  [^6] This result is not limited to square lattices but applies more generally to topologies where qubits can be ordered such that no qubit is directly connected to more than two of its predecessors.. As for the interleaved CB instances, building on the simple single CZ gate example, it can be shown that just two additional instances of interleaved CB —where extra S gates are added before each CZ gate—are required to capture all the necessary additional (products of) fidelities. Consequently, a single layer of parallel CZ gates can be fully characterized using a total of $11$ CB instances. Here, a single instance refers to a CB experiment with a fixed initial Pauli state and measurement basis. For the four layers shown in Fig. 2, this approach requires a total of $44$ CB instances.
This number can be further reduced to $16$ instances (four per layer) by generalizing the RC procedure to include S gates in the twirling set. This extension symmetrizes the $X$ and $Y$ components of the noise, reducing the amount of information that needs to be extracted and thereby decreasing the experimental runtime [40]. It is important to emphasize that the discussion here focuses solely on the number of unique CB instances required—an essential aspect of implementing the procedure efficiently [40]. However, this does not address the sampling complexity of the characterization strategies, which is also an important factor in determining the overall runtime. A more comprehensive discussion of the sampling complexity associated with Pauli noise characterization can be found in Ref. [41].
IV Multi-layer CB
In realistic scenarios, quantum circuits typically involve more than a single (dressed) Clifford layer, requiring the characterization of an entire set $\mathcal{C}$ of distinct layers [15, 8]. The conventional approach to this problem involves characterizing each unique layer individually, following the procedure outlined earlier. In this section, we propose augmenting the conventional approach with the novel MLCB strategy that leverages circuits retaining the fundamental CB structure but relying on building blocks that consist of combinations of different layers, picked from the set $\mathcal{C}$. A simple-yet-illustrative example is shown in Fig. 3(a), where a combination of two layers, depicted in blue and green, is repeated $d$ times. In the remainder of this Section, we demonstrate the potential of this technique by focusing at first on the characterization of a simple system and then discussing its generalization to more complicated and relevant scenarios.
Figure 3: (a) Illustration of the MLCB protocol. The structure is analogous to the one shown in Fig. 1 with the important difference that, here, the “building block” which is repeated $d$ times is not a single Clifford layer but a combination of more layers (in this specific case two, shown in blue and green). (b) Simple scenario where MLCB allows to measure with high accuracy the product $f^{B}_{IXI}f^{G}_{XZX}$ of two fidelities, belonging to two different layers, which cannot be accessed with conventional CB protocols. (c) Simple system consisting of three qubits and two layers.
IV.1 Simple example
The simple setup shown in Fig. 3(c) consists of three qubits arranged in a line and numbered from 00 to $2$. The goal is to characterize the two depicted layers, which correspond to a single CZ between qubits 00 and $1$ (blue layer) and between qubits $1$ and $2$ (green layer). The conventional approach, described in the previous Section, prescribes to focus on each layer individually and perform CB protocols to determine with high accuracy all the (products of) fidelities related to the orbits that involve the Pauli strings $\alpha\in\mathcal{K}$. Since each layer consists of a single CZ gate, there are four unlearnable DOF, which can be identified with the set of weight-one fidelities
$\left\{f^{B}_{XII},f^{B}_{IXI},f^{G}_{IXI},f^{G}_{IIX}\right\}.$ (11)
Here we use the notation $f^{L}_{\alpha}$ for the Pauli fidelities, where $\alpha$ is a Pauli string (ordered according to the increasing qubit indexes) and $L\in\{B,G\}$ indicates the associated layer, i.e. the blue (B) one or the green (G) one. To complete the characterization task, all the fidelities in (11) must be estimated using low-accuracy methods, such as symmetry assumptions or unit-depth variants of CB protocols.
Remarkably, the MLCB experiment shown in Fig. 3(b) enhances the characterization process by enabling the high-accuracy characterization of one additional DOF. To demonstrate this, it is crucial to determine the behavior of a Pauli operator $P_{\alpha}$ under conjugation by the alternating sequence of blue and green layers, which constitutes the core of MLCB circuits. The first $m\geq 0$ layers of this sequence, starting from the leftmost blue layer, are associated with the unitary
$\bar{U}^{(m)}=\begin{cases}(U_{B}U_{G})^{m/2}&\text{ if }m\text{ is even },\\ U_{B}(U_{G}U_{B})^{(m-1)/2}&\text{ if }m\text{ is odd },\end{cases}$ (12)
where $U_{B}$ and $U_{G}$ are the unitary associated with the blue and green layer, respectively. Let us define the sequence of Pauli strings $\alpha^{(m)}$ by requiring that
$P_{\alpha^{(m)}}=\bar{U}^{(m)}P_{\alpha}[\bar{U}^{(m)}]^{\dagger}.$ (13)
In the specific example of Fig. 3(b), with the initial string $\alpha=\alpha^{(0)}=XIX$, the sequence is $2$-periodic: $\alpha^{(2k)}=XIX$ and $\alpha^{(2k+1)}=XZX$ for $k\geq 0$. If we prepare a $+1$ eigenstate of operator $P_{XIX}$, implement $d$ pairs of the blue and green layer and then measure (with readout twirling) the expectation value $\tilde{\mathsf{O}}_{d}^{XIX}$ of the same operator, we obtain
$\tilde{\mathsf{O}}_{d}^{XIX}=A_{XIX}(f^{B}_{XIX}f^{G}_{XZX})^{d}.$ (14)
In analogy with the standard CB protocols, the contribution of SPAM errors is encapsulated in the prefactor $A_{XIX}$ and a simple exponential fit of data obtained for different $d$ provide us with an high-accuracy estimation of the product $o_{3}=f^{B}_{XIX}f^{G}_{XZX}$.
The two fidelities in $o_{3}$ differ form the ones in Eq. 11 and, if we were considering a general Pauli noise model, they would therefore not help in imposing constraints on the fidelities in Eq. 11. Importantly, however, in the context of effective noise models, which feature a limited number of parameters and thus allow for scalable characterization, not all the Pauli fidelities are independent quantities. Specifically, considering the SPL model, we show in Appendix B that
$f^{B}_{XII}=\frac{f^{B}_{XIX}}{f^{B}_{IIX}}$ (15)
and
$\begin{split}f^{G}_{IIX}=\frac{f^{G}_{XZI}(f^{G}_{IZX}f^{G}_{IIX})}{f^{G}_{XZX}f^{G}_{IZI}},\end{split}$ (16)
which leads to
$\mu_{1}^{BG}\equiv\frac{f^{B}_{XII}}{f^{G}_{IIX}}=o_{3}\,\frac{f^{G}_{IZI}}{f^{B}_{IIX}f^{G}_{XZI}(f^{G}_{IZX}f^{G}_{IIX})}.$ (17)
The meaning of the notation $\mu_{1}^{BG}$ will become apparent later. All the terms on the RHS can be determined with high accuracy, using MLCB for $o_{3}$ and standard CB protocols for all the other terms. Consequently, MLCB grants us access to an additional DOF, which can be identified with the fidelity ratio $\mu_{1}^{BG}$, which was unlearnable under the conventional approach alone.
To extend this result to more complex scenarios, it is crucial to formalize the concept of equivalence between functions of fidelities. Specifically, we define two functions as equivalent if one can be fully determined from the other using only high-accuracy data, i.e. (products of) learnable fidelities. Within this framework, Eq. (17) demonstrates that $\mu_{1}^{BG}$ and $o_{3}$ are equivalent, providing a foundation for applying this principle in more intricate cases.
IV.2 Generalization to two arbitrary layers
Figure 4: (a) An example of several disconnected graphs $Gi$ obtained by considering edges associated with two layers, blue and green, of parallel two-qubit gates. The graphs can be divided into open and closed chains, as shown by the black rectangles. (b) Analysis of open chains showing the conjugations of specific Pauli strings $\alpha^{(m)}$ (with $m=0,\dots 4$) under four subsequent layers, in the pattern BGBG, that constitute the building block of MLCB protocols. Different chains, featuring four and five qubits, are depicted on the left (each with a specific qubit numbering). By looking at the corresponding qubits (and gates) on the right side one can identify the circuit blocks that allow to measure specific products of fidelities. The latter can be then used to determine the ratios $\mu_{q}^{BG}$ of unlearnable fidelities associated with the qubit $q$ (highlighted in yellow). Gray-shaded areas shows scenarios in which that the presence of additional qubit is trivial. (c) Analysis of closed chains consisting of four qubits. As for the previous panel, the circuits shows how MLCB can measure a product of four fidelities, which can then be used to determine the ratio $\mu_{q}^{BG}$ (with $q=0$ highlighted in yellow).
The MLCB scheme can be leveraged to enhance learnability in more general scenarios. In this section, we consider the characterization of two generic $n$-qubit (dressed) layers, identified with blue (B) and green (G) colors. We begin by deriving conclusions applicable to general layers consisting of parallel two-qubit Clifford gates and then proceed by specifically focusing on the relevant case of layers consisting of parallel CZ gates on a QPU with square lattice topology.
A useful way to visualize the problem at hand is to represent each 2Q gate as a colored edge connecting the two qubits in its support. Those edges collectively form several disconnected graphs $G_{i}$ over the QPU. An example of this visualization, applied to a square topology, is shown in Fig. 4(a). Since the Clifford layers consist of parallel 2Q gates, a qubit cannot simultaneously belong to two edges of the same color. Consequentially, each graph $G_{i}$ is a one-dimensional chain, either open or closed, composed of edges with alternating colors. Moreover, closed chains necessarily consist of an even number of qubits. This simple structure, which arises independently of the QPU topology, significantly simplifies the analysis of MLCB protocols, as each graph $G_{i}$ can be analyzed independently and in parallel.
A straightforward generalization of the three-qubit example discussed earlier shows that, by preparing a $+1$ eigenstate of a Pauli operator $P_{\alpha}$ and executing $d$ pairs of blue and green layers, the expectation value $\tilde{\mathsf{O}}_{d}^{\alpha}$ of $P_{\alpha^{(2d)}}$ reads
$\tilde{\mathsf{O}}_{d}^{\alpha}\propto\prod_{j=1}^{d}f^{B}_{\alpha^{(2j-2)}}f^{G}_{\alpha^{(2j-1)}}.$ (18)
The more complex task is determining how to select the Pauli string $\alpha$ (or a set of them) so that the associated sequence $\alpha^{(m)}$, defined by Eq. (13), yields expectation values $\tilde{\mathsf{O}}_{d}^{\alpha}$ that allow us to learn DOF which would have been otherwise unlearnable. A couple of preliminary observations prove to be particularly helpful in this endeavour. As anticipated before, it is possible to focus on a single graph $G_{i}$ at a time, since the single-qubit Pauli operators in $\alpha^{(m)}$ associated with the qubits in $G_{i}$, are fully determined by the single-qubit Pauli operators in $\alpha^{(0)}=\alpha$ associated with the same qubits in $G_{i}$. Moreover, the sequence $\alpha^{(m)}$ is periodic, with a period that depends on the specific Clifford layers. In relevant cases, such as for layers consisting exclusively of CZ gates, the period is particularly short as it cannot exceed $4$, since in this case $(U_{B}U_{G})^{2}=I$.
In the remainder of this Section, we thoroughly analyze all the possible graphs $G_{i}$ that arise when characterizing a pair of layers consisting of parallel CZ gates on a square topology. This allows us to prove that, for every qubit $q$ in the bulk of a graph $G_{i}$, the MLCB protocol enables the high-accuracy determination of an additional DOF that would have been unlearnable using conventional CB methods alone. In particular, this DOF can be identified with the ratio $\mu_{q}^{(BG)}$ of two individual unlearnable fidelities
$\mu_{q}^{(BG)}=\frac{f^{B}_{\beta}}{f^{G}_{\gamma}}.$ (19)
Here the Pauli string $\beta$ ($\gamma$) consists only of identities $I$ with the exception of a single $X$ associated to the qubit connected to $q$ via the blue (green) edge. This definition explains the usage of the symbol $\mu_{1}^{(BG)}$ in Eq. (17), where qubit number $1$ is the only one in the bulk of the graph shown in Fig. 3(c).
Remarkably, all of those DOF can be accessed by preparing Pauli strings $\alpha$ consisting solely of $I$ and $X$ operators, regardless of the specific qubit $q$ or graph $G_{i}$. Consequently, by preparing a $+1$ eigenstate of $X^{\otimes n}$ and measuring in the $X$ basis, it becomes possible to determine all the fidelity ratios $\mu_{q}^{(BG)}$ simultaneously, with a single MLCB execution. This parallelization introduces only a limited overhead compared to the conventional procedure. Together with the details in Appendix B, the remainder of this section provides a concrete and operational framework for implementing MLCB in relevant scenarios, highlighting the protocol’s capability to unlock additional DOF in an efficient way.
IV.2.1 Open chains
Let us start with the analysis of open chains, some of which are depicted in the left rectangle of Fig. 4(a). Open chains consisting of just two qubits feature a single layer (and have no bulk), meaning that MLCB cannot be implemented. As for open chains with three qubits, they have been already fully covered in Section IV.1. We therefore proceed by focusing on open chains featuring a total of four qubits. To analyze them within the MLCB framework, it is necessary to repeatedly implement the minimal four-layer block with a BGBG pattern, as shown in Fig. 4(b). If the open chain features two blue CZ gates, Fig. 4(b) shows how MLCB can directly measure the product of four fidelities $o_{4}=f^{B}_{XIXI}f^{G}_{XZXZ}f^{B}_{XIXZ}f^{G}_{XZXI}$. In Appendix B, we prove it to be equivalent to the ratio $\mu_{1}^{(BG)}$ between two unlearnable fidelities associated with the bulk qubit $1$. Similarly, if the open chain features two green CZ gates, MLCB can directly measure $o^{\prime}_{4}=f^{B}_{IXIX}f^{G}_{IXZX}f^{B}_{ZXIX}f^{G}_{ZXZX}$, which we prove to be equivalent to the ratio $\mu_{2}^{(BG)}$. By exploiting inversion symmetry, analogous analyses can be performed to determine the ratios of unlearnable fidelities associated with the other qubits in the bulk of four-qubit chains.
As for open chains consisting of five qubits, as shown in Fig. 4(b), there are three qubits in the bulk. The ratios $\mu_{q}^{(BG)}$ associated with qubits $q=1,3$ can be readily determined by following the strategies discussed above for the four-qubit chains, as the relevant Pauli strings $\alpha^{(m)}$ always have an identity $I$ operator on the qubit at the opposite side of the chain (see the gray-shaded areas). By contrast, to determine the ratio associated with the central qubit, all the five qubits are involved. In particular, MLCB can directly measure the product $o_{5}=f^{B}_{IXIXI}f^{G}_{IXZXZ}f^{B}_{ZXIXZ}f^{G}_{ZXZXI}$, which we then prove in Appendix to be equivalent to $\mu_{2}^{(BG)}$. Interestingly, even if more qubits were present (see the gray-shaded areas), the relevant Pauli strings $\alpha^{(m)}$ always feature operators different from the identity $I$ on no more than five qubits. As a result, when analyzing longer open chains, we can readily determine the ratio $\mu_{q}^{(BG)}$ for each qubit in the bulk by resorting to the MLCB strategies detailed so far.
IV.2.2 Closed chains
The shortest closed chain consists of just two qubits, both residing in the bulk, each paired by a CZ gate in both layers (see in the upper left corner of the “Closed chains” rectangle in Fig. 4(a)). In this straightforward case, the MLCB approach directly provides us with two fidelity products, $c_{2}=f^{B}_{IX}f^{G}_{ZX}$ and $c^{\prime}_{2}=f^{B}_{XI}f^{G}_{XZ}$, which are equivalent to the two ratios $\mu_{0}^{(BG)}$ and $\mu_{1}^{(BG)}$. The analysis of chains consisting of four qubits is more involved. In the case shown in Fig. 4(c), MLCB allows to measure $c_{4}=f^{B}_{IXXX}f^{G}_{ZXYY}f^{B}_{IYYX}f^{G}_{ZYXY}$, a product of fidelities that is equivalent to the ratio $\mu_{q}^{(BG)}$ of unlearnable fidelities associated with qubit $q=0$. By rotating and/or mirroring the chain, one can systematically derive the ratios for the other three qubits $q$ in the bulk of the chain. Detailed derivation is provided in Appendix B. For longer chains involving six or more qubits, the results derived for open chains can be directly applied. This is because the latter never involves more than five qubits at a time, regardless of the chain length. A clear example of this can be seen from Fig. 4(b) by identifying (as a single qubit number $5$) the two gray-shaded qubits, effectively showing how to use MLCB to determine the ratio $\mu_{q}^{(CB)}$ associated with each qubit in a bulk of a six-qubit closed chain.
IV.3 More than two layers
The approach detailed so far can be readily generalized to scenarios featuring more than two layers. Let us focus on a specific qubit $q$, which belongs to the support of $l$ layers in the set $\{L_{1},\dots,L_{l}\}$. By simply implementing the procedure outlined before, for each pair of layers $(L_{i}L_{j})$ we can thus determine the ratio $\mu_{q}^{(L_{i}L_{j})}$, which would have been unlearnable if only standard CB were used. Unfortunately, not all ratios $\mu_{q}^{(L_{i}L_{j})}$ are independent from each other. Specifically, the knowledge of the $l-1$ ratios $\mu_{q}^{(L_{i}L_{i+1})}$ (for $i=1,\dots,l-1$) is enough to reconstruct all the remaining ones, meaning that MLCB provides us with $l-1$ DOF that would have been unlearnable otherwise.
Let us discuss the concrete effect of MLCB on the highly structured example in Fig. 2(a), where every qubit $q$ in the bulk of the QPU belongs to the support of each one of the four layers, blue (B), green (G), red (R) and orange(O). By performing MLCB on three pairs of layers, we can determine three independent ratios per qubit, say $\mu_{q}^{(BG)}$, $\mu_{q}^{(GR)}$, and $\mu_{q}^{(RO)}$. In the limit of large $n$, this brings down the total number of unlearnable DOF that needs to be determined with low-accuracy methods from $\sim 4n$ to $\sim n$. Importantly, this $75\%$ reduction of unlearnable DOF comes with a rather small experimental overhead, around $7\%$ more circuit runs, since it requires to perform only $3$ extra MLCB experiments on top of the $44$ CB instances required in any case by the standard approach. This shows how MLCB can be beneficial, especially in the highly relevant context of characterizing errors during the execution of circuits with few dense layers of CZ gates.
V Experimental implementation
We now demonstrate the practical feasibility of implementing the MLCB protocol on IQM GarnetTM, a 20-qubit superconducting quantum computer, and highlight its ability to uncover mischaracterization stemming from (unjustified) symmetry assumptions on the system fidelities. Specifically, we characterize four Clifford layers, consisting of parallel CZ gates on the 20-qubit chip [22] as shown in Fig. 5(a). Given this layer arrangement, the longest graphs $G_{i}$ associated with layers pairs are squares, like the one shown in Fig. 4(c). As discussed in the Appendix B, square-shaped graphs are among the most challenging to analyze when it comes to connecting products of fidelities directly measured by MLCB to the fidelity ratio $\mu_{q}^{L_{1}L_{2}}$ associated with the same DOF. This complexity arises because several fidelities measured using standard CB approaches must be combined together. The successful implementation of MLCB in such a demanding scenario is therefore indicative of the general effectiveness of the protocol.
Figure 5: (a) Layout of the IQM GarnetTM [22] quantum processor, outlining the four characterized layers. The square graph, analyzed in greater detail, is highlighted with the gray circle. (b) Experimental comparison of fidelity ratios $\mu_{q}^{(BG)}$ derived using MLCB (crosses) with the corresponding low-accuracy estimates $\tilde{\mu}_{q}^{(BG)}$, obtained using conventional CB protocols under symmetry assumption (bullet markers). Results from the first (purple) and second (orange) characterization runs are shown. Violin plots represent the distribution of the bootstrapped data ($200$ resamplings), while error bars indicate the associated standard deviations. The averaged results over the two runs are shown in black. (c) Analysis of the differences $\tilde{\mu}_{q}^{BG}-\tilde{\mu}_{q}^{BG}$, averaged over both runs, for all four qubits. Each characterization run included $8$ circuit depths $d$ for CB and $6$ for MLCB. For each depth and state preparation stage, $10^{4}$ shots were performed, distributed across $30$ RC instances.
Our objective is to compare the high-accuracy fidelity ratios $\mu_{q}^{(L_{1}L_{2})}$ obtained using MLCB with the same quantities computed relying solely on standard CB methods and symmetry assumption, denoted with $\tilde{\mu}_{q}^{(L_{1}L_{2})}$. While performing the experiments, we observed drifts in time of several (products of) fidelities, which complicate the comparison of the fidelity ratios. Indeed, the drifts raise ambiguity as to whether discrepancies between $\mu_{q}^{(L_{1}L_{2})}$ and $\tilde{\mu}_{q}^{(L_{1}L_{2})}$ stem from the poor accuracy of the symmetry assumption or simply from the time-dependent change in fidelities. To mitigate this issue, we performed two back-to-back characterization runs—encompassing both standard CB and MLCB—and ranked different subregions of the QPU according to the temporal stability of their fidelities.
The most stable graph $G_{i}$ corresponds to the square highlighted in Fig. 5(a), with qubits relabeled to align with the scheme in Fig. 4(b). A detailed analysis of the experimental results associated with this graph is provided in Fig. 5(b), where we report the fidelity ratios $\mu_{q}^{BG}$ and $\tilde{\mu}_{q}^{BG}$ for each of the four qubits $q=0,\dots,3$ for each of the two characterization runs: the first shown in purple, the second in orange. To estimate the associated statistical uncertainties, we use bootstrapping, resampling the experimental data $200$ times with respect to both the shots and the RC instances, giving rise to the distribution shown in the violin plots. The associated standard deviations are shown as error bars. For both $\tilde{\mu}_{q}^{BG}$ (bullet markers) and $\mu_{q}^{BG}$ (cross markers), the results from the two runs are statistically consistent within the uncertainties. We therefore proceed by computing the values averaged over the two runs, depicted in black.
[…]
This experimental case demonstrates the capability of MLCB to improve characterization accuracy in practical scenarios. However, it also underscores the challenges posed by time drifts and statistical uncertainties, which can reduce the benefits of enhanced learnability provided by MLCB. In this respect, we observe that the statistical uncertainties associated with $\mu_{q}^{BG}$ are significantly larger than the ones associated with $\tilde{\mu}_{q}^{BG}$. Were this not the case, the higher accuracy enabled by MLCB would have been even more impactful and apparent. The disparity in the uncertainties arises because $\tilde{\mu}_{q}^{BG}$ is derived almost directly from a small subset of CB data, whereas $\mu_{q}^{BG}$ is calculated from the directly measured $c_{4}$ value through an equivalence relation that involves multiple additional CB measurements (as detailed in Tab. 5 in Appendix B).
While this issue warrants further exploration, it is not a critical limitation for several reasons. First, in this comparison, the calculation of $\mu_{q}^{BG}$ was chosen for convenience, enabling a direct comparison with $\tilde{\mu}_{q}^{BG}$. From a learnability perspective, however, the directly measured $c_{4}$—which has a significantly smaller statistical uncertainty—is equally valid and fully exposes the advantages of MLCB. Moreover, the statistical uncertainty of $c_{4}$ could be further reduced by optimizing resource allocation, such as dedicating more experimental shots to MLCB runs. Lastly, as noted earlier, the square graph analyzed here represents one of the most challenging cases in terms of connecting MLCB outcomes with the fidelity ratio $\mu_{q}^{BG}$. For other graph types, this connection is more straightforward and introduces a much smaller increase in statistical uncertainty.
VI Improved error characterization
Having proved that MLCB can significantly reduce the number of unlearnable DOF in realistic scenarios, this raises the question to which extent this additional knowledge can improve the overall accuracy of the fitted SPL model. Note that this is not a trivial step, as it requires combining high-accuracy data, obtained from CB and MLCB, with low-accuracy ones, necessary to make up for the limited learnability of Pauli noise.
We answer this important point by numerically simulating the characterization of a $20$-qubit IQM GarnetTM QPU [22]. In particular, we consider two different dense layer structures, depicted in the two insets of Fig. 6(a). The rightmost one corresponds to the configuration used for the experiment in Sec. V, with pairs of layers defining closed square chains, whose MLCB analysis follow the scheme shown in Fig. 4(c). By contrast, in the leftmost, the relevant chains are open, with variable lengths, and can thus be analyzed according to the schemes in Fig 4(b).
Our numerical simulations consist of five steps. At first, (i) we randomly generate a realistic SPL noise model, i.e. we associate a map $\Lambda_{L}$ to each one of the four layer $L\in\{B,G,R,O\}$ based on model parameters $\vec{\lambda}^{L}$ that mimics the behavior of realistic devices, see Appendix D for more details. (ii) From each exact noise map $\Lambda_{L}$ we compute all the (products of) fidelities that are needed to fit the SPL model. In an actual experiment, those would be obtained using conventional CB, MLCB, as well as possible alternative low-accuracy methods. (iii) We artificially add uncertainties independently to all (product of) fidelities, in the form of Gaussian noise, to simulate the statistical uncertainty associated with an actual experimental characterization. For (products of) fidelities determined via conventional CB and MLCB, we consider a fixed standard deviation $\sigma$. To simulate the implementation of low-accuracy strategies that involves performing extra measurements, like the one based on unit-depth CB-like circuits, we take into account for the additional uncertainties due to SPAM noise and the lack of multiplicative precision by considering a larger standard deviation $\sigma^{\prime}\geq\sigma$. (iv) We then use these noisy fidelities to fit two SPL models for each layer: $\Lambda_{L}^{c}$, which is based on parameters $\{\vec{\lambda}^{L}_{c}\}$ fitted using solely conventional CB data and low-accuracy strategies, and $\Lambda_{L}^{m}$, whose fitted parameters $\{\vec{\lambda}^{L}_{m}\}$ also leverages the additional knowledge associated with MLCB experiments. (v) Finally, we compute the $L_{1}$ distances $\Delta_{t}$, with $t=c,m$, between the exact and the reconstructed noise models, summed over all the layers $L\in\{B,G,R,O\}$ as
$\Delta_{t}=\sum_{L}||\vec{\lambda}^{L}-\vec{\lambda}^{L}_{t}||_{1}.$ (20)
We also define their ratio as $r=\Delta_{m}/\Delta_{c}$.
Figure 6: (a) Reduction of the $L_{1}$ distance between the exact and characterized noise model enabled by MLCB. We consider $200$ randomly generated noise models and plot the ratio $r=\Delta_{m}/\Delta_{c}$, where the conventional strategy $c$ relies on the symmetry assumption, for each of the two layer configurations shown in the insets. (b) Similarly to the previous panel, we plot ratios $r$ associated with conventional strategies that allow to measure unlearnable quantities with accuracy $\sigma^{\prime}\geq\sigma=10^{-4}$, for the configuration shown in left inset of panel (a).
The most intricate aspect of this otherwise straightforward procedure lies in step (iv), particularly the role of fidelity products measured with MLCB during the model fitting process. Several strategies can be adopted to address this step. One straightforward method is to treat experimental data from conventional CB, MLCB, and low-accuracy methods on the same ground, by generalizing Eq. (8) to the multi-layer scenario
$\log(\vec{\xi}^{+})=-2\bar{M}^{+}\vec{\lambda}^{+}.$ (21)
Here, $\vec{\lambda}^{+}$ is the concatenation of the parameter vectors ${\vec{\lambda}}_{L}$, associated with each individual layer $L\in\{B,G,R,O\}$. Similarly, $\vec{\xi}^{+}$ is the concatenation of the (product of) fidelities vectors $\vec{\xi}_{L}$, which are used in the conventional approach, supplemented with the products of fidelities measured using MLCB. Eq. 21 can be then solved using non-negative least square fitting in a $4|\mathcal{K}|$-dimensional space.
Alternatively, for each layer $L$, one can first identify a specific set of unlearnable fidelities $\{f^{L}_{\beta}\}$ required to fit the SPL model. For example, as discussed above, one can focus on all weight-one Pauli strings $\beta$ consisting of a single $X$ operator on the layer’s support. Since these fidelities are inherently unlearnable, their initial estimates rely on low-accuracy data. However, they can be significantly improved by imposing the high-accuracy constraints provided by MLCB, e.g., $\mu_{q}^{(L_{i}L_{j})}$. This optimization step, detailed in Appendix C, yields refined fidelity estimates $\{f^{L}_{\beta}\}$, which can then be used for fitting the SPL model layer-by-layer. In the reminder of this work, we primarily adopt this second approach, which places greater emphasis on leveraging high-accuracy MLCB data over low-accuracy estimates. However, other fitting strategies are certainly possible. A thorough exploration of such strategies represents a compelling avenue for future research, as it could unlock further improvements in model fitting accuracy and performance.
The numerical results are summarized in Fig.6, where we plot the values of the ratio $r$ for different scenarios. In panel (a) we visualize the values of $r$ for $200$ randomly generated noise model, considering both the layer structures shown in the two insets. We consider a reasonable experimental uncertainty of $\sigma=10^{-4}$ [8] and we use the symmetry assumption to estimate the unlearnable individual fidelities. Regardless of the specific layer structure, we observe a $r$ consistently below $1$, showing that the possibility to measure with high accuracy extra DOF, unlocked by MLCB, leads to a better overall error characterization. In Fig. 6(b), we simulate a characterization procedure that rely on extra measurements, for example via unit-depth CB-like experiments, to estimate with low-accuracy the required unlearnable fidelities. For a fixed $\sigma=10^{-4}$, we consider different uncertainties associated with unlearnable quantities $\sigma\leq\sigma^{\prime}\leq 100\sigma$ and compute the ratio $r$ for $100$ randomly generated noise models for each value of $\sigma^{\prime}$. Provided that $\sigma^{\prime}\gtrsim 3\sigma$, which makes the distinction between high- and low-accuracy procedures meaningful and justified, we observe values of the ratio $r$ well below $1$, which saturates to a range $0.5\lesssim 0.8$ in the realistic regime $\sigma^{\prime}\gtrsim 10\sigma$. This shows again how MLCB can significantly improve the error characterization. By contrast, in the extreme and unrealistic regime $\sigma\sim\sigma^{\prime}$, MLCB can even be disadvantageous due to the amplification of the $\sigma$ error associated with the computation of ratios $\mu_{1}^{(L_{1},L_{j})}$ (discussed in the previous Section and detailed in Appendix B) that makes it more convenient to just consider the direct measurements of the unlearnable fidelities.
VII Effects on error mitigation
Error characterization plays a crucial role in informing noise-aware error mitigation strategies, prompting an important question: how does MLCB and the associated improved characterization impact error mitigation performance? To address this, we conducted a series of numerical experiments analyzing the performance of PEC [13, 8, 15] on Clifford circuits. The choice of Clifford circuits is particularly advantageous because evaluating PEC performance in this context is both conceptually straightforward and computationally efficient [23]. Moreover, generic non-Clifford circuits can always be decomposed into a sum of maps associated with Clifford circuits, making the following analysis relevant, at least at a qualitative level, also in practical and general use cases. More precisely, data obtained on Clifford circuits can be indicative of worst-case scenarios performances [23].
In the analysis we focus on circuits consisting of the four (dressed) Clifford layers depicted in the rightmost inset of Fig. 6(a), each one associated with a randomly generated noise map $\Lambda_{L}$. As discussed in the previous section, for each layer $L\in\{B,G,Y,R\}$ we also define a low-accuracy reconstructed noise map $\Lambda_{L}^{c}$ as well as another reconstructed map $\Lambda_{C}^{m}$ that makes use of the extra high-accuracy DOF provided by MLCB. When using PEC to mitigate errors on a given layer, we probabilistically modify the circuit such that the noisy execution of the layer is effectively composed with the inverse of the characterized noise map $\Lambda_{L}^{t}$, with $t\in\{c,m\}$. As a result, the execution of each layer is effectively associated with the map $\xi_{L}^{(t)}=\Phi(U_{L})\circ\Lambda_{L}\circ(\Lambda_{L}^{t})^{-1}$, which corresponds to the noiseless layer execution in the limit of perfect characterization, i.e. $\Lambda_{L}^{t}=\Lambda_{L}$. The effect of such a map on a Pauli operator $P_{\alpha}$ simply reads
$\xi_{L}^{(t)}[P_{\alpha}]=P_{\alpha^{\prime}}R_{\alpha,t}^{L}$ (22)
where $P_{\alpha^{\prime}}=\Phi(U_{c})[P_{\alpha}]$ is the conjugation of $P_{\alpha}$ under the Clifford unitary and
$R_{\alpha,t}^{L}=\frac{f_{\alpha}^{L}}{f_{\alpha,t}^{L}}$ (23)
is the ratio between the exact Pauli fidelity $f^{L}_{\alpha}$, derived from $\Lambda_{L}$, and the fidelity $f_{\alpha,t}^{L}$ obtained from the characterized model $\Lambda_{L}^{t}$, with $t\in\{c,m\}$.
We consider Clifford circuits, consisting of the repeated implementations of $J$ Clifford layers $\bar{L}_{j}$, with $j\in[1,J]$. Each $\bar{L}_{i}$ is the concatenation of a layer randomly sampled from $\{B,G,R,O\}$ with a random layer consisting of 1Q Clifford gates. By preparing a $+1$ eigenstate of a Pauli operator $P_{\beta^{(0)}}$ and measuring the expectation value $O$ of its conjugation $P_{\beta^{(J)}}$ with respect to all Clifford layers, we obtain (neglecting SPAM errors and applying PEC)
$O_{t}=\prod_{i=0}^{J-1}R_{\beta^{(i)},t}^{L}.$ (24)
Here $\beta^{(i)}$ is the Pauli string obtained under conjugation by the first $i\geq 0$ Clifford layers, i.e. $P_{\beta^{(i)}}=\Phi(\prod_{j=1}^{i}U_{\bar{L}_{j}})[P_{\beta^{(0)}}]$. A perfect characterization would result in the noiseless result $O=1$ and deviation from this value are consequence of the finite accuracy of the reconstructed maps $\Lambda_{L}^{t}$.
Figure 7: Bias $1-O_{t}$ resulting from the application of PEC on random Clifford circuits under $200$ random SPL noise models, comparing two characterization methods: the conventional procedure ($t=c$) and the advanced one leveraging the additional learnability provided by MLCB ($t=m$). The MLCB-based method achieves improved error mitigation accuracy, consistently outperforming the standard procedure regardless of the observable’s weight, whether low ($W=2$) or high ($W=20$).
By computing $O_{c}$ and $O_{m}$ for different noise models and different circuits, we observed that the latter features a much smaller standard deviation, indicating that the extra learnability offered by MLCB can directly result into a better performance of PEC. The results of this numerical analysis are shown in Fig. 7, where we consider the same $200$ randomly generated noise models employed for Fig. 6(a). For each noise model, we calculated $O_{t}$ across ten random circuits, with each circuit composed of 40 random Clifford layers $\bar{L}_{j}$. Additionally, each circuit began with a randomly chosen initial Pauli string $\beta^{(0)}$ , subject only to the condition that the final measured Pauli string $\beta^{(J)}$ had a specified weight $W$. The latter parameter does not seem to play a significant role. The standard deviation associated with $O_{c}$ is $\text{std}(O_{c})=0.08$ for both $W=2$ and $W=20$. By contrast, the standard deviation associated with $O_{m}$ is $\text{std}(O_{m})=0.03$ for both $W=2$ and $W=20$, showing almost a three-fold improvement over the conventional scenario.
VIII Discussion
In this paper, we have introduced the MLCB protocol and showed its potential in increasing the accuracy of characterization of effective Pauli noise models, thanks to an improved learnability. Numerical simulations and a simple experiment have been performed focusing on practically relevant scenarios, featuring dense Clifford layers consisting of parallel CZ gates implemented on a QPU with square topology. In these cases, we show in detail how to set up the MLCB procedure to significantly decrease the number of unlearnable DOF while keeping the runtime overhead at a minimum. Our results showcase a scalable and resource-efficient procedure that can be readily extended to various combinations of CZ layers with minimal overhead.
While our focus has been on CZ gates and (SPL) models with weight two generators, MLCB can be generalized to other Clifford gates, higher-weight SPL models or alternative effective Pauli noise frameworks like CER [16]. These extensions, though valuable, are beyond the scope of this paper. On the other hand, MLCB is naturally compatible with generalizations of RC beyond the Pauli group with which the independent number of noise parameters can be reduced [40], as exemplified in Section II.2 for CZ gates and considering the addition of $S$ gates to the twirling set.
It is also important to recognize that limited learnability is not the sole factor affecting characterization accuracy and the performance of noise-aware error mitigation: time-dependent parameter drift, as observed in our experiment, and out-of-model noise contributions can also play a significant role. Recent work, such as Ref. [23], has explored these two issues in the context of PEC. In this sense, our discussion in SectionVI complements these analyses by focusing on the effect of limited learnability. In general, it is important to stress that practical implementations must consider all these factors holistically to ensure robust characterization and effective noise mitigation.
The MLCB implementation presented in this work is both practical and resource-efficient, but there is room for optimization in several areas.
First, we have not addressed the optimal allocation of experimental resources, such as the distribution of shots and rounds of RC, between different CB and MLCB protocols. In this regard, insights from the statistical analyses in protocols like ACES, as detailed in [12], provide valuable directions for future work. Clearly, optimizing the resource allocation could enhance the precision of characterization within a finite experimental runtime. More broadly, the study of sampling complexity associated with quantum noise characterization is an active and compelling area of research, with recent advancements detailed, e.g., in [42, 43, 41].
Second, as discussed in Section VI, efficiently fitting noise model parameters based on measured (products of) fidelities is a non-trivial task. While we propose a weighting strategy that prioritizes high-accuracy data (obtained through CB and MLCB) over lower-accuracy data from other methods, this approach, validated through numerical simulations of random noise models, is not guaranteed to be optimal. Further exploration in this area could yield more robust and efficient fitting techniques, enhancing the overall utility of MLCB.
Finally, while our implementation significantly reduces the number of unlearnable fidelities with only a minor runtime increase, we do not exclude the possibility of alternative strategies that could unlock even more DOF, potentially at the cost of increased experimental runtime. In this respect, during the finalization of this work, we became aware of theoretical advancements in [44], which propose a general framework for computing the learnability of Pauli noise models, considering also the SPAM parameters as target of learning. Our results are consistent with their findings in overlapping cases, such as the characterization of quasi-local noise on $n$-qubit systems arranged in 1D rings with two dense layers of parallel CZ gates, thus suggesting the optimality of MLCB in this scenario. In general, we believe our work complements the theoretical framework in [44] by providing practical details for characterizing general layers on 2D QPU s with square topologies, thus enabling straightforward implementation on physical devices, and by quantifying the impact of MLCB’s enhanced learnability on characterization quality and its effectiveness in error mitigation within the widely adopted SPL model.
IX Acknowledgment
A.C. gratefully acknowledges S. Chen and A. Kandala for their insightful discussions during the APS March Meeting 2023, where the foundational ideas of this work have been presented. The authors acknowledge support from the German Federal Ministry of Education and Research (BMBF) under Q-Exa (grant No. 13N16062) and QSolid (grant No. 13N16161). The authors also acknowledge the entire IQM Technology team for their support in the development of this work.
Appendix A The example of a single CZ gate
Let us focus on a system with $n=2$ qubits and on a layer $C$ consisting of a single CZ gate between the two qubits, a simple-yet-instructive example to discuss the limited learnability of Pauli noise. The fifteen non-trivial Pauli operators $P_{\alpha}\in\mathbb{P}_{2}$ are listed in the first row of Tab. 1 in the main text, while the second row shows their conjugation under $C$. Since there are only three orbits with cardinality one, associated with the Pauli strings $\alpha\in\{IZ,ZI,ZZ\}$, the three corresponding Pauli fidelities are the only ones that can be determined with high accuracy using standard Cycle Benchmarking (CB). See, for example, the setup in Fig. 8(a) that can measure $A_{ZZ}(f_{ZZ})^{d}$ and, thus, determine with high accuracy the individual fidelity $f_{ZZ}$. All other orbits have cardinality $2$ (which is the largest possible since, in this case, $U_{C}^{2}=I$). In several cases, however, the Pauli operators belonging to the same orbit share the same two-qubit support (for example, $\mathcal{R}^{C}_{XX}=\mathcal{R}^{C}_{YY}=\{P_{XX},P_{YY}\}$). In these cases, the orbit’s cardinality can be reduced to one by using interleaved CB and, in particular, by inserting two extra $S$ gates after each execution of the layer $C$ (and then recompiling them with the following layer of single-qubit (1Q) gates). This modifies the associated unitary as $U_{\bar{C}}=(S\otimes S)U_{C}$, providing us with the third row of Tab. 1 and, importantly, with four novel orbits with size one: $\mathcal{R}^{\bar{C}}_{\alpha}$ for $\alpha\in\{XX,XY,YX,YY\}$. An example of an interleaved CB circuit that allows to determine $f_{XX}$ is shown in Fig. 8(b). At this point, we are left with eight Pauli strings, which can be conveniently organized into two sets $\mathcal{S}_{A}=\{IX,IY,ZX,ZY\}$ and $\mathcal{S}_{B}=\{XI,YI,XZ,YZ\}$, whose associated orbits cannot be reduced to size one. This means that the Pauli fidelities corresponding to the elements of each set therefore cannot be determined individually with high accuracy. However, it is important to stress that, out of the $8$ DOF associated with these $8$ fidelities, the CB protocols still allows us to determine $6$ independent high-accuracy constraints on the products of fidelity pairs, thus leaving us with only $2$ unlearnable DOF. Indeed, focusing on set $\mathcal{S}_{A}$, one can determine the product $f_{IX}f_{ZX}$, as shown in Fig. 8(c), as well as $f_{IY}f_{ZY}$ and $f_{IX}f_{ZY}$ (the fourth product is not independent since $f_{IY}f_{ZX}=(f_{IY}f_{ZY})(f_{IX}f_{ZX})/(f_{IX}f_{ZY})$). The same applies to the other set $\mathcal{S}_{B}$. All in all, this example shows that a single CZ gate is associated with $2$ unlearnable DOF, which cannot be determined using CB protocols and which can be (non-unambiguously) identified with fidelities $f_{IX}$ and $f_{XI}$. These results mirror the findings discussed in Ref. [19] for a single CNOT gate.
Figure 8: (a) Simple example of a CB circuit that allows for the high-accuracy determination of Pauli fidelity $f_{ZZ}$. For the sake of clarity, here we do not display the white boxes associated with RC and, instead, we show (in red) the Pauli operator which is measured at the end and how it propagates back (considering its conjugation under the CZ gates) all the way to the state preparation stage. (b) Similarly to the previous panel, we show how interleaved CB can determine the fidelity $f_{XX}$ with high accuracy. The extra purple diamonds show $S$ gates that are applied after every CZ (and eventually recompiled together with the non-displayed layer of 1Q gates) which map the Pauli operators $P_{YY}$ and $P_{XX}$ into each other. (c) In this case, CB protocols can only determine with high accuracy the product of fidelities $f_{XI}f_{XZ}$, which cannot be individually measured with high accuracy.
Appendix B Equivalence between products of fidelities
The objective of this section is to establish the equivalence between products of fidelities directly measurable via Multi-Layer Cycle Benchmarking (MLCB)—such as $o_{3},o_{4},o^{\prime}_{4},o_{5},c_{4}$ introduced in the main text—and specific ratios of unlearnable fidelities, $\mu_{q}^{(BG)}$. As established at the end of Sec. IV.1, two functions of fidelities are deemed equivalent if one can be fully derived from the other using only high-accuracy data, i.e. (products of) learnable fidelities.
To achieve this, we first introduce an equivalence test to formally prove the equivalence between functions of fidelities within a single layer. We then extend this test to multi-layer scenarios and demonstrate its application using a concrete example of three-qubit chains. This example illustrates each step required to prove equivalence and establish relationships such as those expressed in Eqs. (15)–(17). Building on this example, we generalize the methodology to more complex cases and summarize the results.
B.1 Equivalence test
We begin by focusing on a single layer $C$ and linear combinations of logarithms of fidelities. These combinations can be written as $\zeta_{Y}=\sum_{\alpha\in\mathcal{X}_{Y}}\gamma_{\alpha}\log f^{C}_{\alpha}$, where $\gamma^{Y}_{\alpha}$ are numerical coefficients and $\mathcal{X}_{Y}$ is a set of Pauli strings.
Within the Sparse Pauli-Lindblad (SPL) model, the vector of model parameters $\vec{\lambda}$ determines all functions of fidelities via Eq. (8). In particular, the vector $\vec{\Xi}$, consisting of all learnable (products of) fidelities which can be determined, for instance, by following the strategy outlined in Section II.2, is determined by
$\log(\vec{\Xi})=\bar{M}\vec{\lambda}.$ (25)
Analogously, $\zeta_{Y}$ can be determined from the model parameters via
$\zeta_{Y}=(\vec{m}_{Y})^{T}\vec{\lambda}$ (26)
where the $\kappa$-th element of the single-row matrix $(\vec{m}_{Y})^{T}$ reads
$(\vec{m}_{Y})_{\kappa}=\sum_{\alpha\in\mathcal{X}_{Y}}\gamma_{\alpha}\langle\alpha,\kappa\rangle_{sp}.$ (27)
To test whether two functions of fidelities, $\zeta_{A}$ and $\zeta_{B}$, are equivalent, we compute $\bar{M}$, $\vec{m}_{A}$ and $\vec{m}_{B}$. Let $r$ be the rank of $\bar{M}$; $r_{Y}$ the rank of matrix $M$ concatenated with the extra row $(\vec{m}_{Y})^{T}$; $r_{AB}$ the rank of the matrix $\bar{M}$ concatenated with $(\vec{m}_{A})^{T}$ and $(\vec{m}_{B})^{T}$. The equivalence condition is then
$\zeta_{A}\text{ and }\zeta_{B}\text{ are equivalent}\Leftrightarrow r_{A}=r_{B}=r_{AB}.$ (28)
Indeed, $r_{A}=r_{AB}$ shows that $(\vec{m}_{B})^{T}$ is a linear combination of the rows of $\bar{M}$ and of $(\vec{m}_{A})^{T}$, meaning that $\zeta_{B}$ can be expressed in terms of $\zeta_{A}$ and some learnable quantities in $\vec{\Xi}$. The same logic applies symmetrically for $A$ and $B$.
Notably, if $r_{Y}=r$, then $\zeta_{Y}$ is fully learnable and provides no new information for fitting the model parameters.
Figure 9: Visualization of single-row matrices $(\vec{m}_{\alpha})^{T}$ associated with the blue layer (panel a) and the green layer (panel b) for the three-qubit open chain depicted in Fig. 3(c). On the horizontal axes, all the $27$ ordered Pauli strings $\kappa\in\mathcal{K}$ are listed. The first nine consists are weight-one Pauli strings, the second nine are weight-two Pauli strings with support on qubits 00 and $1$, and the last nine are weight-two Pauli strings with support on qubits $1$ and $2$. As for the Pauli strings $\alpha$, they are specified to the left of each row. Note that the last row actually depicts $(\vec{m}_{j}^{T})=(\vec{m}_{IIX}^{T})+(\vec{m}_{IZX}^{T})$.
B.2 Three-qubit open chain
We now consider the three-qubit open chain described in Sec. IV.1. Our goals are: (i) to prove the equivalence between $o_{3}=f^{B}_{XIX}f^{G}_{XZX}$ and $\mu_{1}^{(BG)}=f^{B}_{XII}/f^{G}_{IIX}$ (ii) to prove Eqs. (15-17) which directly link these quantities.
For concreteness, we assume a Quantum Processing Unit (QPU) with square topology, ensuring qubits 00 and $2$ are not neighbours. In this case, the SPL model describing the noise in the chain features a total of $|\mathcal{K}|=27$ parameters, associated with the Pauli strings $\kappa$ listed on the $x$-axes of the plots in Fig. 9.
We first analyze the blue layer, focusing on the function of (now trivial) fidelities $F^{B}_{1}=f^{B}_{XII}$ and $F^{B}_{2}=f^{B}_{XIX}$. Computing the corresponding matrix rows
$(\vec{m}_{XII}^{T})_{\kappa}=\langle XII,\kappa\rangle_{sp}$ (29)
$(\vec{m}_{XIX}^{T})_{\kappa}=\langle XIX,\kappa\rangle_{sp}$ (30)
we verify the equivalence condition from Eq. (28). This proves $\log(F^{B}_{1})$ can be expressed as a linear combination of $\log(F^{B}_{2})$ and other function of learnable fidelities. To explicitly identify the latter, we implement a simple search algorithm, which progressively attempt to get rid of (randomly chosen) rows of matrix $\bar{M}$ while still preserving the condition on the ranks in Eq. (28). We find that, in this case, the single row of matrix $\bar{M}$ associated with the learnable fidelity $f^{B}_{IIX}$, i.e.
$(\vec{m}_{IIX}^{T})_{\kappa}=\langle IIX,\kappa\rangle_{sp},$ (31)
is sufficient to relate $(\vec{m}_{XII}^{T})$ and $(\vec{m}_{XIX}^{T})$ via $(\vec{m}_{XII}^{T})=(\vec{m}_{XIX}^{T})-(\vec{m}_{IIX}^{T})$. This is visualized in Fig. 9(a). In view of the subsequent analysis, it is convenient to summarize this result as
$\log(F^{L}_{1})=\epsilon^{L}\log(F^{L}_{2})+(\vec{\Sigma}^{L})^{T}\log(\vec{F}^{L})$ (32)
with $L=B$, $\epsilon^{B}=1$, $\vec{\Sigma}^{B}=(-1)$ and $\vec{F}^{B}=(f^{B}_{IIX})$. This proves the validity of 15.
Similarly, for the green layer we focus on functions $F^{G}_{1}=f^{G}_{IIX}$ and $F^{G}_{2}=f^{G}_{XZX}$, proven to be equivalent by (28). Moreover, we show that the corresponding matrix rows $(\vec{m}_{IIX}^{T})$ and $(\vec{m}_{XZX}^{T})$ can be related to each other by using three rows of matrix $\bar{M}$. Specifically, we obtained
$(\vec{m}_{IIX}^{T})=-(\vec{m}_{XZX}^{T})+(\vec{m}_{XZI}^{T})-(\vec{m}_{IZI}^{T})+(\vec{m}_{j}^{T}).$ (33)
where the last row $(\vec{m}_{j}^{T})=(\vec{m}_{IIX}^{T})+(\vec{m}_{IZX}^{T})$ is associated to the learnable product of fidelities in the orbit $\mathcal{R}^{G}_{IIX}=\{P_{IIX},P_{IZX}\}$. Such a relation is visualized in Fig. 9(b). This result can be summarized in the form of Eq. (32), with $L=G$, $\epsilon^{G}=-1$, $\vec{\Sigma}^{G}=(1,-1,1)$ and $\vec{F}^{G}=(f^{G}_{XZI},f^{G}_{IZI},f^{G}_{IIX}f^{G}_{IZX})$. This proves the validity of 16.
By combining these results, valid for each layer individually, we obtain Eq. 17, thus showing the equivalence between $o_{3}=F^{B}_{2}F^{G}_{2}$ and $\mu_{1}^{(BG)}=F^{B}_{1}/F^{G}_{1}$. It is important to stress that the equivalence class does not consists only of these two functions of fidelities. It is indeed possible to find other equivalent functions of fidelities, all associated to the same Degrees of Freedom (DOF), whose learnability is unlocked by MLCB. For example, one has
$f^{B}_{XII}f^{G}_{IZX}=\mu_{1}^{(BG)}(f^{G}_{IIX}f^{G}_{IZX}),$ (34)
where $f^{G}_{IIX}f^{G}_{IZX}$ is learnable with conventional CB.
B.3 Other chains
Longer chains introduce additional challenges due to the connectivity of the underlying QPU. Non-neighboring qubits in the chain may still interact via the hardware’s topology, increasing the number of SPL parameters and altering the relationships between fidelity functions. Relevant examples are shown in the first column of Tables 2-5
In the following, we analyze all relevant chain configurations, considering a QPU with square topology. For each case, we apply the methodology outlined for the three-qubit chain. Specifically, focusing on one layer at a time, we use the equivalence test to demonstrate the equivalence between two functions of fidelities, denoted as $F^{L}_{1}$ and $F^{L}_{2}$, respectively. This approach enables us to prove the equivalence—and derive a direct connection—between the fidelity product measurable via MLCB, $F^{B}_{2}F^{G}_{2}$, and the ratios of unlearnable fidelities, $F^{B}_{1}/F^{G}_{1}$. The results for all chain configurations are summarized in Tables 2-5.
It is important to note that the reported $\vec{\Sigma}^{L}$ and $\vec{F}^{L}$ are not unique, as alternative choices for these quantities can still satisfy Eq. (32). An interesting open question concerns the optimization of the choice of $\vec{\Sigma}^{L}$ and $\vec{F}^{L}$, with the goal of minimizing the number of components in each vector for a given chain. While this optimization does not affect the learnability of fidelities and is thus beyond the scope of this paper, it can be beneficial when implementing the characterization procedure in practice. For instance, in the four-qubit closed chain analyzed in Table 5, connecting $c_{4}=f^{B}_{IXXX}f^{G}_{ZXYY}f^{B}_{IYYX}f^{G}_{ZYXY}$ to the simple fidelity ratio $\mu_{q}^{(BG)}$ requires considering a total of $28$ learnable (products of) fidelities. Reducing this number could streamline the analyses and make them more robust with respect to the spread of statistical uncertainties.
Appendix C Fitting the noise parameters
Here we detail how we impose the high-accuracy constraint obtained with MLCB on the low-accuracy fidelities. Let us focus on a qubit $q$, belonging to the support of $l$ layers $L_{1},\dots,L_{l}$. When relying exclusively on conventional CB protocols, qubit $q$ is associated with $l$ unlearnable DOF, which can be identified with the $l$ fidelities in $\{f^{L_{j}}_{\beta_{q}^{j}}\}$, where the weight-one Pauli strings $\beta_{q}^{j}$ feature a single $X$ operator on the qubit connected to $q$ via the $L_{j}$ layer. Low-accuracy methods, like the symmetry assumption or unit-depth CB-like experiments, provide us with estimates for each one of these fidelities, which we indicate with an extra tilde $\tilde{f}^{L_{j}}_{\beta_{q}^{j}}$ and represent the numerical values used for the conventional model fitting.
MLCB allows us to determine $l-1$ additional high-accuracy constraints, which can be identified, for example, with fidelity ratios
$\mu_{q}^{L_{j},L_{j+1}}=\frac{f^{L_{j}}_{\beta_{q}^{j}}}{f^{L_{j+1}}_{\beta_{q}^{j+1}}}$ (35)
for $j=1,\dots,l-1$. This reduces the number of unlearnable DOF to one, which we can identify with a specific fidelity, say $f_{q}^{L_{1}}$. We estimate its numerical value $u$ by minimizing the sum $S$ of the squared residues
$S=(\tilde{f}^{L_{j}}_{\beta_{q}^{j}}-u)^{2}+\sum_{j=1}^{l-1}\left(\tilde{f}^{L_{j}}_{\beta_{q}^{j}}-u\prod_{i=1}^{j}\mu_{q}^{L_{i},L_{i+1}}\right)^{2}$ (36)
computed by comparing the low-accuracy fidelity estimations with the effect of the high-fidelity constraints from MLCB. This procedure can be readily adjusted to tackle scenarios where different pairs of layers have been studied with MLCB.
Appendix D Random noise models
The random noise models are generated to simulate the behavior of realistic devices. The procedure involves assigning values to all model parameters for a given layer $L$ distinguishing them based on the following criteria:
  • Weight ($w$) of the Pauli operator $P_{\kappa}$, either $w=1$ or $w=2$;
  • Status ($s$) of the involved qubits, which can be either active ($s=a$) if all qubits in the support of $P_{\kappa}$ belong to the support of $L$, or inactive ($s=i$) otherwise.
Inactive parameters are drawn from a Gaussian distribution with mean $m_{w}^{i}$ and standard deviation $\sigma_{w}^{i}$. The active ones are sampled from a Gaussian with mean $m_{w}^{a}$ and standard deviation $\sigma_{w}^{a}$, where, for every gate in the layer, $m_{w}^{a}$ itself is drawn from a Gaussian with mean $\bar{m}_{w}^{a}$ and standard deviation $\bar{\sigma}_{w}^{a}$ This additional sampling step introduces variability between gates within a single layer, resulting in better and worse-performing ones.
Finally, note that any negative model parameters resulting from the sampling process are set to zero. The parameter values used in this procedure reads
[…]
for $w=1$ and
[…]
for $w=2$.
Table 2: Four-qubit open chains (blue-green-blue)
[Uncaptioned image] $F^{B}_{1}$ $F_{2}^{B}$ $\epsilon^{B}$ $\vec{\Sigma}^{B}$ $\vec{F}^{B}$
$f^{B}_{XIII}$ $f^{B}_{XIXI}f^{B}_{XIXZ}$ $+0.50$ $\begin{pmatrix}-0.50\\ \end{pmatrix}$ $\begin{pmatrix}f^{B}_{IIXZ}f^{B}_{IIXI}\\ \end{pmatrix}$
$F^{G}_{1}$ $F_{2}^{G}$ $\epsilon^{G}$ $\vec{\Sigma}^{G}$ $\vec{F}^{G}$
$f^{G}_{IIXI}$ $f^{G}_{XZXZ}f^{G}_{XZXI}$ $-0.50$ $\begin{pmatrix}+0.50\\ -0.50\\ +0.25\\ +0.75\\ \end{pmatrix}$ $\begin{pmatrix}f^{G}_{XZII}f^{G}_{XZII}\\ f^{G}_{IZII}f^{G}_{IZII}\\ f^{G}_{IIXZ}f^{G}_{IZXZ}\\ f^{G}_{IZXI}f^{G}_{IIXI}\\ \end{pmatrix}$
[Uncaptioned image] $F^{B}_{1}$ $F_{2}^{B}$ $\epsilon^{B}$ $\vec{\Sigma}^{B}$ $\vec{F}^{B}$
$f^{B}_{XIII}$ $f^{B}_{XIXI}f^{B}_{XIXZ}$ $+0.50$ $\begin{pmatrix}+0.25\\ -0.50\\ +0.25\\ -0.25\\ \end{pmatrix}$ $\begin{pmatrix}f^{B}_{IIIZ}f^{B}_{IIIZ}\\ f^{B}_{IIXZ}f^{B}_{IIXI}\\ f^{B}_{XZII}f^{B}_{XIII}\\ f^{B}_{XIIZ}f^{B}_{XZIZ}\\ \end{pmatrix}$
$F^{G}_{1}$ $F_{2}^{G}$ $\epsilon^{G}$ $\vec{\Sigma}^{G}$ $\vec{F}^{G}$
$f^{G}_{IIXI}$ $f^{G}_{XZXZ}f^{G}_{XZXI}$ $-0.50$ $\begin{pmatrix}-0.50\\ +0.25\\ +0.75\\ -0.25\\ +0.50\\ -0.25\\ +0.25\\ \end{pmatrix}$ $\begin{pmatrix}f^{G}_{IZII}f^{G}_{IZII}\\ f^{G}_{IIXZ}f^{G}_{IZXZ}\\ f^{G}_{IIXI}f^{G}_{IZXI}\\ f^{G}_{IIIZ}f^{G}_{IIIZ}\\ f^{G}_{XZII}f^{G}_{XZII}\\ f^{G}_{XIII}f^{G}_{XIII}\\ f^{G}_{XIIZ}f^{G}_{XIIZ}\\ \end{pmatrix}$
Table 3: Four-qubit open chains (green-blue-greem
[Uncaptioned image] $F^{B}_{1}$ $F_{2}^{B}$ $\epsilon^{B}$ $\vec{\Sigma}^{B}$ $\vec{F}^{B}$
$f^{B}_{IXII}$ $f^{B}_{IXIX}f^{B}_{ZXIX}$ $+0.50$ $\begin{pmatrix}-0.25\\ +0.25\\ -0.50\\ \end{pmatrix}$ $\begin{pmatrix}f^{B}_{ZXII}f^{B}_{ZXZI}\\ f^{B}_{IXZI}f^{B}_{IXII}\\ f^{B}_{IIIX}f^{B}_{IIIX}\\ \end{pmatrix}$
$F^{G}_{1}$ $F_{2}^{G}$ $\epsilon^{G}$ $\vec{\Sigma}^{G}$ $\vec{F}^{G}$
$f^{G}_{IIXI}$ $f^{G}_{XZXZ}f^{G}_{XZXI}$ $-0.50$ $\begin{pmatrix}-0.50\\ +0.50\\ +1.00\\ \end{pmatrix}$ $\begin{pmatrix}f^{G}_{IIZI}f^{G}_{IIZI}\\ f^{G}_{IXZI}f^{G}_{ZXZI}\\ f^{G}_{IIZX}f^{G}_{IIIX}\\ \end{pmatrix}$
[Uncaptioned image] $F^{B}_{1}$ $F_{2}^{B}$ $\epsilon^{B}$ $\vec{\Sigma}^{B}$ $\vec{F}^{B}$
$f^{B}_{IXII}$ $f^{B}_{IXIX}f^{B}_{ZXIX}$ $+0.50$ $\begin{pmatrix}+0.25\\ -0.25\\ -0.25\\ +0.25\\ -0.25\\ \end{pmatrix}$ $\begin{pmatrix}f^{B}_{ZIII}f^{B}_{ZIII}\\ f^{B}_{IIIX}f^{B}_{IIIX}\\ f^{B}_{ZXII}f^{B}_{ZXZI}\\ f^{B}_{IXZI}f^{B}_{IXII}\\ f^{B}_{ZIIX}f^{B}_{ZIIX}\\ \end{pmatrix}$
$F^{G}_{1}$ $F_{2}^{G}$ $\epsilon^{G}$ $\vec{\Sigma}^{G}$ $\vec{F}^{G}$
$f^{G}_{IIXI}$ $f^{G}_{XZXZ}f^{G}_{XZXI}$ $-0.50$ $\begin{pmatrix}-0.25\\ +0.25\\ +0.50\\ -0.50\\ +0.75\\ \end{pmatrix}$ $\begin{pmatrix}f^{G}_{ZIII}f^{G}_{ZIII}\\ f^{G}_{ZIIX}f^{G}_{ZIZX}\\ f^{G}_{IXZI}f^{G}_{ZXZI}\\ f^{G}_{IIZI}f^{G}_{IIZI}\\ f^{G}_{IIZX}f^{G}_{IIIX}\\ \end{pmatrix}$
Table 4: Five-qubit open chains
[Uncaptioned image] $F^{B}_{1}$ $F_{2}^{B}$ $\epsilon^{B}$ $\vec{\Sigma}^{B}$ $\vec{F}^{B}$
$f^{B}_{IXIII}$ $f^{B}_{IXIXI}f^{B}_{ZXIXZ}$ $+0.50$ $\begin{pmatrix}-0.25\\ +0.25\\ -0.50\\ \end{pmatrix}$ $\begin{pmatrix}f^{B}_{ZXIII}f^{B}_{ZXZII}\\ f^{B}_{IXIII}f^{B}_{IXZII}\\ f^{B}_{IIIXI}f^{B}_{IIIXZ}\\ \end{pmatrix}$
$F^{G}_{1}$ $F_{2}^{G}$ $\epsilon^{G}$ $\vec{\Sigma}^{G}$ $\vec{F}^{G}$
$f^{G}_{IIIXI}$ $f^{G}_{IXZXZ}f^{G}_{ZXZXI}$ $-0.50$ $\begin{pmatrix}-0.50\\ +0.75\\ +0.25\\ +0.50\\ \end{pmatrix}$ $\begin{pmatrix}f^{G}_{IIZII}f^{G}_{IIZII}\\ f^{G}_{IIZXI}f^{G}_{IIIXI}\\ f^{G}_{IIIXZ}f^{G}_{IIZXZ}\\ f^{G}_{IXZII}f^{G}_{ZXZII}\\ \end{pmatrix}$
[Uncaptioned image] $F^{B}_{1}$ $F_{2}^{B}$ $\epsilon^{B}$ $\vec{\Sigma}^{B}$ $\vec{F}^{B}$
$f^{B}_{IXIII}$ $f^{B}_{IXIXI}f^{B}_{ZXIXZ}$ $+0.50$ $\begin{pmatrix}+0.25\\ -0.25\\ -0.25\\ +0.25\\ -0.25\\ \end{pmatrix}$ $\begin{pmatrix}f^{B}_{IXZII}f^{B}_{IXIII}\\ f^{B}_{ZIIXI}f^{B}_{ZIIXZ}\\ f^{B}_{ZXIII}f^{B}_{ZXZII}\\ f^{B}_{ZIIII}f^{B}_{ZIIII}\\ f^{B}_{IIIXZ}f^{B}_{IIIXI}\\ \end{pmatrix}$
$F^{G}_{1}$ $F_{2}^{G}$ $\epsilon^{G}$ $\vec{\Sigma}^{G}$ $\vec{F}^{G}$
$f^{G}_{IIIXI}$ $f^{G}_{IXZXZ}f^{G}_{ZXZXI}$ $-0.50$ $\begin{pmatrix}+0.25\\ +0.25\\ -0.50\\ +0.50\\ -0.25\\ +0.50\\ \end{pmatrix}$ $\begin{pmatrix}f^{G}_{IIIXZ}f^{G}_{IIZXZ}\\ f^{G}_{ZIIXI}f^{G}_{ZIZXI}\\ f^{G}_{IIZII}f^{G}_{IIZII}\\ f^{G}_{IXZII}f^{G}_{ZXZII}\\ f^{G}_{ZIIII}f^{G}_{ZIIII}\\ f^{G}_{IIIXI}f^{G}_{IIZXI}\\ \end{pmatrix}$
[Uncaptioned image] $F^{B}_{1}$ $F_{2}^{B}$ $\epsilon^{B}$ $\vec{\Sigma}^{B}$ $\vec{F}^{B}$
$f^{B}_{IXIII}$ $f^{B}_{IXIXI}f^{B}_{ZXIXZ}$ $+0.50$ $\begin{pmatrix}+0.50\\ -0.50\\ -0.25\\ -0.25\\ +0.25\\ \end{pmatrix}$ $\begin{pmatrix}f^{B}_{IXZII}f^{B}_{IXIII}\\ f^{B}_{IIIXI}f^{B}_{IIIXZ}\\ f^{B}_{IXIIZ}f^{B}_{IXZIZ}\\ f^{B}_{ZXIII}f^{B}_{ZXZII}\\ f^{B}_{IIIIZ}f^{B}_{IIIIZ}\\ \end{pmatrix}$
$F^{G}_{1}$ $F_{2}^{G}$ $\epsilon^{G}$ $\vec{\Sigma}^{G}$ $\vec{F}^{G}$
$f^{G}_{IIIXI}$ $f^{G}_{IXZXZ}f^{G}_{ZXZXI}$ $-0.50$ $\begin{pmatrix}+0.25\\ +0.50\\ -0.25\\ +0.25\\ +0.75\\ -0.25\\ -0.50\\ \end{pmatrix}$ $\begin{pmatrix}f^{G}_{IXIIZ}f^{G}_{ZXIIZ}\\ f^{G}_{IXZII}f^{G}_{ZXZII}\\ f^{G}_{ZXIII}f^{G}_{IXIII}\\ f^{G}_{IIIXZ}f^{G}_{IIZXZ}\\ f^{G}_{IIIXI}f^{G}_{IIZXI}\\ f^{G}_{IIIIZ}f^{G}_{IIIIZ}\\ f^{G}_{IIZII}f^{G}_{IIZII}\\ \end{pmatrix}$
Table 5: Four-qubit closed chain
[Uncaptioned image] $F^{B}_{1}$ $F_{2}^{B}$ $\epsilon^{B}$ $\vec{\Sigma}^{B}$ $\vec{F}^{B}$
$f^{B}_{IXII}$ $f^{B}_{IXXX}f^{B}_{IYYX}$ $+0.50$ $\begin{pmatrix}+0.25\\ -0.50\\ -0.25\\ -0.25\\ -0.25\\ -0.25\\ -0.25\\ +0.50\\ +0.25\\ +0.25\\ +0.50\\ -0.25\\ \end{pmatrix}$ $\begin{pmatrix}f^{B}_{IIXY}f^{B}_{IIXY}\\ f^{B}_{IIXY}f^{B}_{IIYX}\\ f^{B}_{IXIX}f^{B}_{ZXZX}\\ f^{B}_{ZIII}f^{B}_{ZIII}\\ f^{B}_{IIXX}f^{B}_{IIXX}\\ f^{B}_{IYIX}f^{B}_{ZYZX}\\ f^{B}_{IIZI}f^{B}_{IIZI}\\ f^{B}_{IIZX}f^{B}_{IIIX}\\ f^{B}_{ZIZI}f^{B}_{ZIZI}\\ f^{B}_{IXII}f^{B}_{ZXII}\\ f^{B}_{IXII}f^{B}_{ZYII}\\ f^{B}_{ZYII}f^{B}_{IYII}\\ \end{pmatrix}$
$F^{G}_{1}$ $F_{2}^{G}$ $\epsilon^{G}$ $\vec{\Sigma}^{G}$ $\vec{F}^{G}$
$f^{G}_{IIXI}$ $f^{G}_{ZXYY}f^{G}_{ZYXY}$ $-0.50$ $\begin{pmatrix}-0.25\\ +0.25\\ +0.25\\ -0.25\\ +0.25\\ +0.25\\ +0.50\\ -0.25\\ -0.25\\ +0.25\\ +0.50\\ -0.25\\ +0.25\\ -0.50\\ -0.25\\ +0.25\\ \end{pmatrix}$ $\begin{pmatrix}f^{G}_{ZZII}f^{G}_{ZZII}\\ f^{G}_{IXIY}f^{G}_{IXIY}\\ f^{G}_{IIXY}f^{G}_{ZZXY}\\ f^{G}_{IIYI}f^{G}_{ZIYI}\\ f^{G}_{ZXII}f^{G}_{ZXIZ}\\ f^{G}_{ZYII}f^{G}_{ZYIZ}\\ f^{G}_{IXIX}f^{G}_{IYIY}\\ f^{G}_{IXIZ}f^{G}_{IXII}\\ f^{G}_{ZIII}f^{G}_{ZIII}\\ f^{G}_{IZII}f^{G}_{IZII}\\ f^{G}_{IIXI}f^{G}_{ZIYI}\\ f^{G}_{IXIX}f^{G}_{IXIX}\\ f^{G}_{ZIXI}f^{G}_{IIXI}\\ f^{G}_{IIIY}f^{G}_{IZIY}\\ f^{G}_{IYII}f^{G}_{IYIZ}\\ f^{G}_{IIYY}f^{G}_{ZZYY}\\ \end{pmatrix}$